# PSformer: Parameter-efficient Transformer with Segment Attention for Time Series Forecasting

### v1

@@ B.2 Full Results
Models PSformer SAMformer TSMixer PatchTST MOMENT ModernTCN FEDformer GPT4TS Autoformer RLinear iTransformer Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 96 0.352 0.385 0.381 0.402 0.388 0.408 0.414 0.419 0.387 0.410 0.373 0.399 0.376 0.415 0.376 0.397 0.435 0.446 0.386 0.395 0.386 0.405 192 0.385 0.406 0.409 0.418 0.421 0.426 0.460 0.445 0.410 0.426 0.407 0.419 0.423 0.446 0.416 0.418 0.456 0.457 0.437 0.424 0.441 0.436 336 0.411 0.424 0.423 0.425 0.430 0.434 0.501 0.466 0.422 0.437 0.436 0.437 0.444 0.462 0.442 0.433 0.486 0.487 0.479 0.446 0.487 0.458 720 0.440 0.456 0.427 0.449 0.440 0.459 0.500 0.488 0.454 0.472 0.467 0.474 0.469 0.492 0.477 0.515 0.515 0.517 0.481 0.470 0.503 0.491 Avg 0.397 0.418 0.410 0.424 0.420 0.431 0.468 0.455 0.418 0.436 0.421 0.432 0.428 0.454 0.428 0.426 0.473 0.477 0.446 0.434 0.454 0.448 ETTh2 96 0.272 0.337 0.295 0.358 0.305 0.367 0.302 0.348 0.288 0.345 0.271 0.339 0.332 0.374 0.285 0.342 0.332 0.368 0.288 0.338 0.297 0.349 192 0.335 0.379 0.340 0.386 0.350 0.393 0.388 0.400 0.349 0.386 0.332 0.382 0.407 0.446 0.354 0.389 0.426 0.434 0.374 0.390 0.380 0.400 336 0.356 0.411 0.350 0.395 0.360 0.404 0.426 0.433 0.369 0.408 0.365 0.411 0.400 0.447 0.373 0.407 0.477 0.479 0.415 0.426 0.428 0.432 720 0.389 0.431 0.391 0.428 0.402 0.435 0.431 0.446 0.403 0.439 0.402 0.441 0.412 0.469 0.406 0.441 0.453 0.490 0.420 0.440 0.427 0.445 Avg 0.338 0.390 0.344 0.391 0.354 0.400 0.387 0.407 0.352 0.394 0.343 0.393 0.388 0.434 0.355 0.395 0.422 0.443 0.374 0.398 0.383 0.407 ETTm1 96 0.282 0.336 0.329 0.363 0.327 0.363 0.329 0.367 0.293 0.349 0.310 0.356 0.326 0.390 0.292 0.346 0.510 0.492 0.355 0.376 0.334 0.368 192 0.321 0.360 0.353 0.378 0.356 0.381 0.367 0.385 0.326 0.368 0.340 0.373 0.365 0.415 0.332 0.372 0.514 0.495 0.391 0.392 0.377 0.391 336 0.352 0.380 0.382 0.394 0.387 0.397 0.399 0.410 0.352 0.384 0.373 0.392 0.392 0.425 0.366 0.394 0.510 0.492 0.424 0.415 0.426 0.420 720 0.413 0.412 0.429 0.418 0.441 0.425 0.454 0.439 0.405 0.416 0.420 0.418 0.446 0.458 0.417 0.421 0.527 0.492 0.487 0.450 0.491 0.459 Avg 0.342 0.372 0.373 0.388 0.378 0.392 0.387 0.400 0.344 0.379 0.361 0.384 0.382 0.422 0.351 0.383 0.515 0.493 0.414 0.407 0.407 0.410 ETTm2 96 0.167 0.258 0.181 0.274 0.190 0.284 0.175 0.259 0.170 0.260 0.168 0.261 0.180 0.271 0.173 0.262 0.205 0.293 0.182 0.265 0.180 0.264 192 0.219 0.292 0.233 0.306 0.250 0.320 0.241 0.302 0.227 0.297 0.231 0.305 0.252 0.318 0.229 0.301 0.278 0.336 0.246 0.304 0.250 0.309 336 0.269 0.325 0.285 0.338 0.301 0.350 0.305 0.343 0.275 0.328 0.272 0.328 0.324 0.364 0.286 0.341 0.343 0.379 0.307 0.342 0.311 0.348 720 0.347 0.376 0.375 0.390 0.389 0.402 0.402 0.400 0.363 0.387 0.375 0.394 0.410 0.420 0.378 0.401 0.414 0.419 0.407 0.398 0.412 0.407 Avg 0.251 0.313 0.269 0.327 0.283 0.339 0.281 0.326 0.259 0.318 0.262 0.322 0.292 0.343 0.267 0.326 0.310 0.357 0.286 0.327 0.288 0.332 Weather 96 0.149 0.200 0.197 0.249 0.189 0.242 0.177 0.218 0.154 0.209 0.154 0.209 0.238 0.314 0.162 0.212 0.249 0.329 0.192 0.232 0.174 0.214 192 0.193 0.243 0.235 0.277 0.228 0.272 0.225 0.259 0.197 0.248 0.207 0.257 0.275 0.329 0.204 0.248 0.325 0.370 0.240 0.271 0.221 0.254 336 0.245 0.282 0.276 0.304 0.271 0.299 0.278 0.297 0.246 0.285 0.248 0.289 0.339 0.377 0.254 0.286 0.351 0.391 0.292 0.307 0.278 0.296 720 0.314 0.332 0.334 0.342 0.331 0.341 0.354 0.348 0.315 0.336 0.337 0.342 0.389 0.409 0.326 0.337 0.415 0.426 0.364 0.353 0.358 0.347 Avg 0.225 0.264 0.261 0.293 0.255 0.289 0.259 0.281 0.228 0.270 0.237 0.274 0.310 0.357 0.237 0.271 0.335 0.379 0.272 0.291 0.258 0.278 Electricity 96 0.133 0.229 0.155 0.252 0.171 0.273 0.181 0.270 0.136 0.233 0.133 0.228 0.186 0.302 0.139 0.238 0.196 0.313 0.201 0.281 0.148 0.240 192 0.149 0.242 0.168 0.263 0.191 0.292 0.188 0.274 0.152 0.247 0.147 0.241 0.197 0.311 0.153 0.251 0.211 0.324 0.201 0.283 0.162 0.253 336 0.164 0.258 0.183 0.277 0.198 0.297 0.204 0.293 0.167 0.264 0.163 0.260 0.213 0.328 0.169 0.266 0.214 0.327 0.215 0.298 0.178 0.269 720 0.203 0.291 0.219 0.306 0.230 0.321 0.246 0.324 0.205 0.295 0.194 0.289 0.233 0.344 0.206 0.297 0.236 0.342 0.257 0.331 0.225 0.317 Avg 0.162 0.255 0.181 0.275 0.198 0.296 0.205 0.290 0.165 0.260 0.160 0.255 0.207 0.321 0.167 0.263 0.214 0.327 0.219 0.298 0.178 0.270 Exchange 96 0.091 0.211 0.161 0.306 0.233 0.363 0.088 0.205 0.098 0.224 0.207 0.342 0.139 0.276 0.091 0.212 0.197 0.323 0.093 0.217 0.086 0.206 192 0.197 0.316 0.246 0.371 0.342 0.437 0.176 0.299 0.201 0.323 0.337 0.437 0.256 0.369 0.183 0.304 0.300 0.369 0.184 0.307 0.177 0.299 336 0.345 0.428 0.368 0.453 0.474 0.515 0.301 0.397 0.387 0.454 0.520 0.553 0.426 0.464 0.328 0.417 0.509 0.524 0.351 0.432 0.331 0.417 720 1.036 0.773 1.003 0.750 1.078 0.777 0.901 0.714 1.062 0.783 1.154 0.810 1.090 0.800 0.880 0.704 1.447 0.941 0.886 0.714 0.847 0.691 Avg 0.417 0.432 0.445 0.470 0.532 0.523 0.367 0.404 0.437 0.446 0.555 0.536 0.478 0.477 0.371 0.409 0.613 0.539 0.378 0.417 0.360 0.350 Traffic 96 0.367 0.257 0.407 0.292 0.409 0.300 0.462 0.295 0.391 0.282 0.391 0.271 0.576 0.359 0.388 0.282 0.597 0.371 0.649 0.389 0.395 0.268 192 0.390 0.272 0.415 0.294 0.433 0.317 0.466 0.296 0.404 0.287 0.403 0.275 0.610 0.380 0.407 0.290 0.607 0.382 0.601 0.366 0.417 0.276 336 0.404 0.274 0.421 0.292 0.424 0.299 0.482 0.304 0.414 0.292 0.410 0.280 0.608 0.375 0.412 0.294 0.623 0.387 0.609 0.369 0.433 0.283 720 0.439 0.294 0.456 0.311 0.488 0.344 0.514 0.322 0.450 0.310 0.451 0.305 0.621 0.375 0.450 0.312 0.639 0.395 0.647 0.387 0.467 0.302 Avg 0.400 0.274 0.425 0.297 0.439 0.315 0.481 0.304 0.415 0.293 0.414 0.283 0.604 0.372 0.414 0.295 0.617 0.384 0.626 0.378 0.428 0.282 $1^{st}$ Count 20 22 2 3 0 0 2 3 2 0 6 3 0 0 0 0 0 0 0 0 2 2

### v2

@@ B.2 Full Results
Models PSformer SAMformer TSMixer PatchTST MOMENT ModernTCN FEDformer GPT4TS Autoformer RLinear iTransformer Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 96 0.352 0.385 0.381 0.402 0.388 0.408 0.414 0.419 0.387 0.410 0.373 0.399 0.376 0.415 0.376 0.397 0.435 0.446 0.386 0.395 0.386 0.405 192 0.385 0.406 0.409 0.418 0.421 0.426 0.460 0.445 0.410 0.426 0.407 0.419 0.423 0.446 0.416 0.418 0.456 0.457 0.437 0.424 0.441 0.436 336 0.411 0.424 0.423 0.425 0.430 0.434 0.501 0.466 0.422 0.437 0.436 0.437 0.444 0.462 0.442 0.433 0.486 0.487 0.479 0.446 0.487 0.458 720 0.440 0.456 0.427 0.449 0.440 0.459 0.500 0.488 0.454 0.472 0.467 0.474 0.469 0.492 0.477 0.515 0.515 0.517 0.481 0.470 0.503 0.491 Avg 0.397 0.418 0.410 0.424 0.420 0.431 0.468 0.455 0.418 0.436 0.421 0.432 0.428 0.454 0.428 0.426 0.473 0.477 0.446 0.434 0.454 0.448 ETTh2 96 0.272 0.337 0.295 0.358 0.305 0.367 0.302 0.348 0.288 0.345 0.271 0.339 0.332 0.374 0.285 0.342 0.332 0.368 0.288 0.338 0.297 0.349 192 0.335 0.379 0.340 0.386 0.350 0.393 0.388 0.400 0.349 0.386 0.332 0.382 0.407 0.446 0.354 0.389 0.426 0.434 0.374 0.390 0.380 0.400 336 0.356 0.411 0.350 0.395 0.360 0.404 0.426 0.433 0.369 0.408 0.365 0.411 0.400 0.447 0.373 0.407 0.477 0.479 0.415 0.426 0.428 0.432 720 0.389 0.431 0.391 0.428 0.402 0.435 0.431 0.446 0.403 0.439 0.402 0.441 0.412 0.469 0.406 0.441 0.453 0.490 0.420 0.440 0.427 0.445 Avg 0.338 0.390 0.344 0.391 0.354 0.400 0.387 0.407 0.352 0.394 0.343 0.393 0.388 0.434 0.355 0.395 0.422 0.443 0.374 0.398 0.383 0.407 ETTm1 96 0.282 0.336 0.329 0.363 0.327 0.363 0.329 0.367 0.293 0.349 0.310 0.356 0.326 0.390 0.292 0.346 0.510 0.492 0.355 0.376 0.334 0.368 192 0.321 0.360 0.353 0.378 0.356 0.381 0.367 0.385 0.326 0.368 0.340 0.373 0.365 0.415 0.332 0.372 0.514 0.495 0.391 0.392 0.377 0.391 336 0.352 0.380 0.382 0.394 0.387 0.397 0.399 0.410 0.352 0.384 0.373 0.392 0.392 0.425 0.366 0.394 0.510 0.492 0.424 0.415 0.426 0.420 720 0.413 0.412 0.429 0.418 0.441 0.425 0.454 0.439 0.405 0.416 0.420 0.418 0.446 0.458 0.417 0.421 0.527 0.492 0.487 0.450 0.491 0.459 Avg 0.342 0.372 0.373 0.388 0.378 0.392 0.387 0.400 0.344 0.379 0.361 0.384 0.382 0.422 0.351 0.383 0.515 0.493 0.414 0.407 0.407 0.410 ETTm2 96 0.167 0.258 0.181 0.274 0.190 0.284 0.175 0.259 0.170 0.260 0.168 0.261 0.180 0.271 0.173 0.262 0.205 0.293 0.182 0.265 0.180 0.264 192 0.219 0.292 0.233 0.306 0.250 0.320 0.241 0.302 0.227 0.297 0.231 0.305 0.252 0.318 0.229 0.301 0.278 0.336 0.246 0.304 0.250 0.309 336 0.269 0.325 0.285 0.338 0.301 0.350 0.305 0.343 0.275 0.328 0.272 0.328 0.324 0.364 0.286 0.341 0.343 0.379 0.307 0.342 0.311 0.348 720 0.347 0.376 0.375 0.390 0.389 0.402 0.402 0.400 0.363 0.387 0.375 0.394 0.410 0.420 0.378 0.401 0.414 0.419 0.407 0.398 0.412 0.407 Avg 0.251 0.313 0.269 0.327 0.283 0.339 0.281 0.326 0.259 0.318 0.262 0.322 0.292 0.343 0.267 0.326 0.310 0.357 0.286 0.327 0.288 0.332 Weather 96 0.149 0.200 0.197 0.249 0.189 0.242 0.177 0.218 0.154 0.209 0.154 0.209 0.238 0.314 0.162 0.212 0.249 0.329 0.192 0.232 0.174 0.214 192 0.193 0.243 0.235 0.277 0.228 0.272 0.225 0.259 0.197 0.248 0.207 0.257 0.275 0.329 0.204 0.248 0.325 0.370 0.240 0.271 0.221 0.254 336 0.245 0.282 0.276 0.304 0.271 0.299 0.278 0.297 0.246 0.285 0.248 0.289 0.339 0.377 0.254 0.286 0.351 0.391 0.292 0.307 0.278 0.296 720 0.314 0.332 0.334 0.342 0.331 0.341 0.354 0.348 0.315 0.336 0.337 0.342 0.389 0.409 0.326 0.337 0.415 0.426 0.364 0.353 0.358 0.347 Avg 0.225 0.264 0.261 0.293 0.255 0.289 0.259 0.281 0.228 0.270 0.237 0.274 0.310 0.357 0.237 0.271 0.335 0.379 0.272 0.291 0.258 0.278 Electricity 96 0.133 0.229 0.155 0.252 0.171 0.273 0.181 0.270 0.136 0.233 0.133 0.228 0.186 0.302 0.139 0.238 0.196 0.313 0.201 0.281 0.148 0.240 192 0.149 0.242 0.168 0.263 0.191 0.292 0.188 0.274 0.152 0.247 0.147 0.241 0.197 0.311 0.153 0.251 0.211 0.324 0.201 0.283 0.162 0.253 336 0.164 0.258 0.183 0.277 0.198 0.297 0.204 0.293 0.167 0.264 0.163 0.260 0.213 0.328 0.169 0.266 0.214 0.327 0.215 0.298 0.178 0.269 720 0.203 0.291 0.219 0.306 0.230 0.321 0.246 0.324 0.205 0.295 0.194 0.289 0.233 0.344 0.206 0.297 0.236 0.342 0.257 0.331 0.225 0.317 Avg 0.162 0.255 0.181 0.275 0.198 0.296 0.205 0.290 0.165 0.260 0.160 0.255 0.207 0.321 0.167 0.263 0.214 0.327 0.219 0.298 0.178 0.270 Exchange 96 0.081 0.197 0.161 0.306 0.233 0.363 0.088 0.205 0.098 0.224 0.207 0.342 0.139 0.276 0.091 0.212 0.197 0.323 0.093 0.217 0.086 0.206 192 0.179 0.299 0.246 0.371 0.342 0.437 0.176 0.299 0.201 0.323 0.337 0.437 0.256 0.369 0.183 0.304 0.300 0.369 0.184 0.307 0.177 0.299 336 0.328 0.412 0.368 0.453 0.474 0.515 0.301 0.397 0.387 0.454 0.520 0.553 0.426 0.464 0.328 0.417 0.509 0.524 0.351 0.432 0.331 0.417 720 0.842 0.689 1.003 0.750 1.078 0.777 0.901 0.714 1.062 0.783 1.154 0.810 1.090 0.800 0.880 0.704 1.447 0.941 0.886 0.714 0.847 0.691 Avg 0.358 0.399 0.445 0.470 0.532 0.523 0.367 0.404 0.437 0.446 0.555 0.536 0.478 0.477 0.371 0.409 0.613 0.539 0.378 0.417 0.360 0.403 Traffic 96 0.367 0.257 0.407 0.292 0.409 0.300 0.462 0.295 0.391 0.282 0.391 0.271 0.576 0.359 0.388 0.282 0.597 0.371 0.649 0.389 0.395 0.268 192 0.390 0.272 0.415 0.294 0.433 0.317 0.466 0.296 0.404 0.287 0.403 0.275 0.610 0.380 0.407 0.290 0.607 0.382 0.601 0.366 0.417 0.276 336 0.404 0.274 0.421 0.292 0.424 0.299 0.482 0.304 0.414 0.292 0.410 0.280 0.608 0.375 0.412 0.294 0.623 0.387 0.609 0.369 0.433 0.283 720 0.439 0.294 0.456 0.311 0.488 0.344 0.514 0.322 0.450 0.310 0.451 0.305 0.621 0.375 0.450 0.312 0.639 0.395 0.647 0.387 0.467 0.302 Avg 0.400 0.274 0.425 0.297 0.439 0.315 0.481 0.304 0.415 0.293 0.414 0.283 0.604 0.372 0.414 0.295 0.617 0.384 0.626 0.378 0.428 0.282 $1^{st}$ Count 20 22 2 3 0 0 2 3 2 0 6 3 0 0 0 0 0 0 0 0 2 2